# existence1.bib

### v1

@@ 3.2 Retractive Directions of Sets
Definition 3.2.1.
Given a nonempty set $X$, a direction $d\in X_{\infty}$ is said to be retractive direction of $X$ if for any sequence $\{x_{k}\}\subseteq X$ converging in the direction $d$ and for any $\rho>0$, there exists an index $K$ (depending on $\rho$) such that
@@
Note that $\mathcal{R}(X)\subseteq X_{\infty}$ by definition, and $0\in\mathcal{R}(X)$. We next provide an example of a convex set that has a non-retractive direction.
@@
Hence, $d=(0,\lambda)$ is not a retractive direction of $X$ for any $\lambda>0$. Thus, $\mathcal{R}(X)=\{0\}$. ∎
We now give an example of a non-convex set which has no nonzero retractive direction.
@@
Now we highlight some related definitions. Most notable is Definition 2.3.1 in [auslender_teboulle_2002] which defines an asymptotically linear set, as follows: a closed set $C\subseteq\mathbb{R}^{n}$ is said to be asymptotically linear if for every $\rho>0$ and each sequence $\{x_{k}\}\subseteq C$ that satisfies $x_{k}\in C$, $\|x_{k}\|\rightarrow+\infty$ and $x_{k}\|x_{k}\|^{-1}\rightarrow\overline{x}$, there exists an index $K$ such that $x_{k}-\rho\overline{x}\in C$ for all $k\geq K$. Note that the directions $\bar{x}$ involved in this definition have unit norm, and such directions are retractive according to our Definition 3.2.1. Further, since $\bar{x}$ cannot be zero, the set of such directions is a subset of $\mathcal{R}(C)$.
From the definition of an asymptotically linear set we can invoke a key example of a retractive set which is a polyhedral set.
@@
Consider a nonempty polyhedral set $X$. The inclusion $\mathcal{R}(X)\subseteq X_{\infty}$ always holds by Definition 3.2.1. By Proposition 2.3.1 in [auslender_teboulle_2002], an asymptotically polyhedral set $X\subseteq\mathbb{R}^{n}$ is asymptotically linear. Since the simplest case of an asymptotically polyhedral set is a polyhedral set, it follows that $X$ is asymptotically linear, i.e., $\mathcal{R}(X)=X_{\infty}$. ∎
Another related definition is Definition 1 in [bertsekas_tseng_2006], which considers $x_{k}\in C_{k}$ for an infinite sequence of nested sets $\{C_{k}\}\subseteq\mathbb{R}^{n}$ i.e., $C_{k+1}\subseteq C_{k}$ for all $k$. The directions $d$ of interest are obtained in the limit, as follows:
@@
By letting $C=C_{k}$ for all $k$, a direction $d$ is retractive according to Definition 1 in [bertsekas_tseng_2006] if, for any associated sequence $\{x_{k}\}\subseteq C$ with $\|x_{k}\|\rightarrow$ and $x_{k}/\|x_{k}\|\to d/\|d\|$, we have that $x_{k}-d\in C$ for all sufficiently large $k$. According to this definition, given a retractive direction $d$ and its associated sequence $\{x_{k}\}$, for any $\rho>0$, we have that
@@
implying that $\{x_{k}\}$ is also associated sequence for the direction $\rho d$ for any $\rho>0$. Thus, the condition $x_{k}-d\in C$ for all $k$ large enough can be written as $x_{k}-\rho d\in C$ for all large enough $k$, implying that $d$ is a retractive direction according to Definition 1 in [bertsekas_tseng_2006] and, also, according to our Definition 3.2.1 with $t_{k}=\|x_{k}\|$ for all $k$.
@@
If $d$ is retractive according to our Definition 3.2.1, then by letting $\rho=\|d\|$, we conclude the $d$ is also retractive according to Definition 1 in [bertsekas_tseng_2006]. As stated previously, the only direction $d$ for which the equivalence does not hold is $d=0$. That is, $0\in\mathcal{R}(X)$ by Definition 3.2.1, but not by Definition 1 of [bertsekas_tseng_2006].
We now state some general properties of retractive directions of a set.
Proposition 3.2.1.
For a nonempty set $X$, the set $\mathcal{R}(X)$ of retractive directions of $X$ is a nonempty cone.
@@
Let $d\in\mathcal{R}(X)$ and let $\lambda\geq 0$ be arbitrary. Let $\{x_{k}\}\subset X$ and $\{t_{k}\}\subseteq\mathbb{R}$ be such that $t_{k}\rightarrow+\infty$ and $x_{k}\cdot t_{k}^{-1}\rightarrow\lambda d$, as $k\rightarrow\infty$, and let $\rho>0$ be arbitrary. Since $d\in\mathcal{R}(X)$ and $\rho\lambda>0$, there exists $K$ such that $x_{k}-\rho\lambda d\in X$ for all $k\geq K$. Therefore, $\lambda d\in\mathcal{R}(X)$. ∎∎
@@
As seen above, we have that each $d_{k}$ is a retractive direction of $X$. However, the limit $(1,0)$ is not a retractive direction of $X$. To show this, we consider a sequence $\{\bar{x}_{k}\}\subset X$ given by
@@
Hence, $(1,0)$ is not a retractive direction of $X$ and, consequently $\mathcal{R}(X)$ is not closed. ∎
@@
Proposition 3.2.2.
Let $X=\cap_{i=1}^{m}X_{i}$ be a nonempty intersection set of closed sets $X_{i}$, for some $m\geq 2$. If $X_{\infty}=\cap_{i=1}^{m}(X_{i})_{\infty}$, then $\cap_{i=1}^{m}\mathcal{R}(X_{i})\subseteq\mathcal{R}(X).$
@@
implying that
$x_{k}-\rho d\in\cap_{i=1}^{m}X_{i}\qquad\hbox{for all $k\geq K$}.$
Thus, $d\in\mathcal{R}(X)$. ∎∎
When the sets $X_{i}$ in Proposition 3.2.2 are closed and convex, the condition $X_{\infty}=\cap_{i=1}^{m}(X_{i})_{\infty}$ is always satisfied, as seen from Proposition 2.2.1.
3.3 Retractive Directions of Functions
@@
By definition, one can show that $\mathcal{R}(f)$ is a cone and $0\in\mathcal{R}(f)$. Furthermore, for a proper function $f$ and a nonempty lower-level set $L_{\gamma}(f)$, we have that
@@
It turns out that, in general, there is no special relationship between the cone $\mathcal{R}(f)$ and the cone $\mathcal{R}(L_{\gamma}(f))$. The following example illustrates that we can have $\mathcal{R}(L_{\gamma}(f))\subseteq\mathcal{R}(f)$ for a non-convex function.
@@
Let $f(s)=\sqrt{|s|}$. Then, for every $\gamma>0$, the lower-level set $L_{\gamma}(f)$ is nonempty and bounded, implying that $\mathcal{R}(L_{\gamma}(f))=\{0\}.$ However, we have that $({\rm epi}f)_{\infty}=\{(d,w)\mid d\in\mathbb{R},w\geq 0\}$, implying that $f_{\infty}(d)=0$ for all $d\in\mathbb{R}$. Moreover, in this case $\mathcal{R}(f)=\mathbb{R}$, and we have for any $\gamma>0$,
$\mathcal{R}(L_{\gamma}(f))=\{0\}\subset\mathcal{R}(f).$
In the following proposition, we establish some properties of $\mathcal{R}(f)$ for a convex function. Interestingly, in this case $\mathcal{R}(f)\subseteq\mathcal{R}(L_{\gamma}(f))$ for a nonempty lower-level set $L_{\gamma}(f)$ (converse inclusion to that of Example 3.9).
Proposition 3.3.1.
Let $f$ be a proper closed convex function. Then, the following relations hold:
  • (a)
@@
  • (b)
    For any nonempty lower-level set $L_{\gamma}(f)$ we have
    $\mathcal{R}(f)\subseteq\mathcal{R}(L_{\gamma}(f)).$
@@
(a) Let $d\in\mathcal{R}(f)$. Then, for any $\{x_{k}\}$ converging in direction $d$ and any $\rho>0$, there is a large enough $K$ such that
$f(x_{k}-\rho d)\leq f(x_{k})\qquad\hbox{for all }k\geq K.$
Let $y_{k}=x_{k}-\rho d$. Then, $\{y_{k}\}$ also converges in direction $d$, and the preceding relation implies that
@@
By Proposition 2.5.2 of [auslender_teboulle_2002] for a proper closed convex function $f$, we have that
@@
Since $\rho>0$ and $f_{\infty}$ is positively homogeneous by Proposition 2.5.1 of [auslender_teboulle_2002], it follows that
$f_{\infty}(d)\geq 0,$
which, combined with the fact that $\mathcal{R}(f)\subseteq\mathcal{K}(f)$, implies that $f_{\infty}(d)=0$.
(b) Let $L_{\gamma}(f)$ be nonempty. To arrive at a contradiction, assume that there is $d\in\mathcal{R}(f)$ such that $d\notin\mathcal{R}(L_{\gamma}(f))$. Since $\mathcal{R}(f)\subseteq\mathcal{K}(f)$, it follows that $d\in\mathcal{K}(f)$. By Proposition 2.3.1, we have that $(L_{\gamma}(f))_{\infty}=\mathcal{K}(f)$, implying that $d\in(L_{\gamma}(f))_{\infty}$. Since $d\in(L_{\gamma}(f))_{\infty}$ but it is a non-retractive direction of the set $L_{\gamma}(f)$, there is a sequence $\{x_{k}\}\subset L_{\gamma}(f)$ converging in direction $d$ and some $\bar{\rho}>0$ such that
@@
Without loss of generality, we may assume that
$f(x_{k}-\bar{\rho}d)\geq\gamma\qquad\hbox{for all }k,$
for otherwise we would just choose a suitable subsequence of $\{x_{k}\}$. Since $\{x_{k}\}\subset L_{\gamma}(f)$, it follows that
@@
Thus, we have that $\{x_{k}\}$ converges in direction $d$ and the direction $d$ belongs to $\mathcal{K}(f)$, but $f(x_{k}-\bar{\rho}d)\geq f(x_{k})$ for all $k$. Hence, the direction $d$ is not retractive for the function $f$, i.e., $d\not\in\mathcal{R}(f)$, which is a contradiction. ∎∎
The inclusion in Proposition 3.3.1(b) can be strict, as seen in the following example.
@@
It can be seen that $f_{\infty}(d)=\langle c,d\rangle$. By Proposition 3.3.1, if a direction $d$ is retractive, then we must have $\langle c,d\rangle=0$. Thus,
$\mathcal{R}(f)=\{d\mid\langle c,d\rangle=0\}\subset\mathcal{R}(L_{\gamma}(f))\qquad\hbox{ for all $\gamma\in\mathbb{R}$ }.$
@@
Consider the function $f(x_{1},x_{2})=e^{-\sqrt{x_{1}x_{2}}}$ for $x_{1}\geq 0$, $x_{2}\geq 0$. The lower-level set $L_{\gamma}(f)$ is nonempty for all $\gamma>0$. Consider $L_{1}(f)$ and $L_{\gamma}(f)$ with $\gamma>0$. For the set $L_{1}(f)$ we have
$L_{1}(f)=\mathbb{R}^{2}_{+},$
which is a polyhedral set, so we have $\mathcal{R}(L_{1}(f))=\mathbb{R}^{2}_{+}$. For the set $L_{\gamma}(f)$ with $\gamma>0$, we have
@@
Finally, we consider polynomial functions. A polynomial $h:\mathbb{R}^{n}\rightarrow\mathbb{R}$ of order $p$ has the following representation [Belousov_Klatte_2002]:
@@
Using the relation $h(x)=\sum_{i=0}^{p}\phi_{i}(x)$, for all $x\in\mathbb{R}^{n}$, and the alternative characterization of the asymptotic function, as given in Theorem 2.3.1, we have
$h_{\infty}(d)$ $=\liminf_{\begin{subarray}{c}t\rightarrow\infty\\ d^{\prime}\rightarrow d\end{subarray}}\frac{h(td^{\prime})}{t}$
$=\liminf_{\begin{subarray}{c}t\rightarrow\infty\\ d^{\prime}\rightarrow d\end{subarray}}\left(\sum_{i=0}^{p}t^{-1}\phi_{i}(td^{\prime})\right)$
@@
where the last equality follows from the fact that $\phi_{0}/t\to 0$, as $t\to\infty$ and the fact that each $\phi_{i}$ is a continuous function. When $\mu(d)\geq 2$, if $\phi_{\mu(d)}(d)>0$, then $h_{\infty}(d)=+\infty$, while if $\phi_{\mu(d)}(d)<0$, then $h_{\infty}(d)=-\infty$. When $\mu(d)=1$, then we are left with $\phi_{1}(d)$. ∎∎
@@
By Lemma 3.3.1 we have that
@@
Let $d$ be such that $h_{\infty}(d)=0$. Then, for every $i=1,\ldots,p$, we have $\phi_{i}(-d)=(-1)^{i}\phi_{i}(d)$, thus implying that $h_{\infty}(-d)=0$. Hence, by Theorem 2.5.3 in [auslender_teboulle_2002] it follows that
@@
where the last inclusion in the preceding relation holds since every direction $d\in\mathcal{C}(h)$ is retractive. By Proposition 3.3.1(a), we have that $\mathcal{R}(h)\subseteq\{d\mid h_{\infty}(d)=0\}$, which implies that the equality holds throughout in (7). ∎∎

### v3

@@ 3.2 Cone of Retractive Directions of Sets
Definition 3.2.
Given a set $X$, a direction $d\in X_{\infty}$ is said to be retractive direction of $X$ if for any sequence $\{x_{k}\}\subseteq X$ converging in the direction $d$ and for any $\rho>0$, there exists an index $K$ (depending on $\rho$) such that
@@
Note that, by definition, we have $\mathcal{R}(X)\subseteq X_{\infty}$, $0\in\mathcal{R}(X)$ and, for $X$ empty, $\mathcal{R}(X)=X_{\infty}=\{0\}$. We next provide examples of a convex set and a nonconvex set that have no nonzero retractive direction.
@@
Thus, $d=(0,\lambda)$ is not a retractive direction of $X$ for any $\lambda>0$, and $\mathcal{R}(X)=\{0\}$. ∎
@@
Now we highlight some related definitions. Most notable is Definition 2.3.1 in [2] which defines an asymptotically linear set, as follows: a closed set $X\subseteq\mathbb{R}^{n}$ is said to be asymptotically linear if for every $\rho>0$ and each sequence $\{x_{k}\}\subseteq X$ that satisfies $x_{k}\in X$, $\|x_{k}\|\rightarrow+\infty$ and $x_{k}\|x_{k}\|^{-1}\rightarrow\overline{x}$, there exists an index $K$ such that $x_{k}-\rho\overline{x}\in X$ for all $k\geq K$. Note that the directions $\bar{x}$ involved in this definition have unit norm, and such directions are retractive according to our Definition 3.2. Further, since $\bar{x}$ cannot be zero, the set of such directions is a subset of $\mathcal{R}(X)$.
By the definition of an asymptotically linear set, a polyhedral set $X$ is retractive i.e., $\mathcal{R}(X)=X_{\infty}$, as detailed below.
@@
Consider a polyhedral set $X$.
By Proposition 2.3.1 in [2], an asymptotically polyhedral set $X\subseteq\mathbb{R}^{n}$ is asymptotically linear. Since the simplest case of an asymptotically polyhedral set is a polyhedral set, it follows that $X$ is asymptotically linear, i.e., $\mathcal{R}(X)=X_{\infty}$. ∎
Another related definition is Definition 1 in [7], which considers $x_{k}\in C_{k}$ for an infinite sequence of nested sets $\{C_{k}\}\subseteq\mathbb{R}^{n}$ i.e., $C_{k+1}\subseteq C_{k}$ for all $k$. The directions $d$ of interest are obtained in the limit, as follows:
@@
By letting $C_{k}=X$ for all $k$, a direction $d$ is retractive according to Definition 1 in [7] if, for any associated sequence $\{x_{k}\}\subseteq X$ with $\|x_{k}\|\rightarrow$ and $x_{k}/\|x_{k}\|\to d/\|d\|$, we have that $x_{k}-d\in X$ for all sufficiently large $k$. According to this definition, given a retractive direction $d$ and its associated sequence $\{x_{k}\}$, for any $\rho>0$, we have that
@@
implying that $\{x_{k}\}$ is also associated sequence for the direction $\rho d$ for any $\rho>0$. Thus, the condition $x_{k}-d\in X$ for all $k$ large enough can be written as $x_{k}-\rho d\in X$ for all large enough $k$, implying that $d$ is a retractive direction according to Definition 1 in [7] and, also, according to our Definition 3.2 with $t_{k}=\|x_{k}\|$ for all $k$.
@@
If $d$ is retractive according to our Definition 3.2, then by letting $\rho=\|d\|$, we conclude the $d$ is also retractive according to Definition 1 in [7]. As stated previously, the only direction $d$ for which the equivalence does not hold is $d=0$. That is, $0\in\mathcal{R}(X)$ by Definition 3.2, but not by Definition 1 of [7].
We now state some general properties of the set $\mathcal{R}(X)$.
Proposition 3.1.
For a set $X$, the set $\mathcal{R}(X)$ of retractive directions of $X$ is a nonempty cone.
@@
If $X=\emptyset$, then $X_{\infty}=\mathcal{R}(X)=\{0\}.$ Assume now that $X\neq\emptyset$. Let $d\in\mathcal{R}(X)$ and let $\lambda\geq 0$ be arbitrary. Let $\{x_{k}\}\subset X$ and $\{t_{k}\}\subseteq\mathbb{R}$ be such that $t_{k}\rightarrow+\infty$ and $x_{k}\cdot t_{k}^{-1}\rightarrow\lambda d$, as $k\rightarrow\infty$, and let $\rho>0$ be arbitrary. Since $d\in\mathcal{R}(X)$ and $\rho\lambda>0$, there exists $K$ such that $x_{k}-\rho\lambda d\in X$ for all $k\geq K$. Therefore, $\lambda d\in\mathcal{R}(X)$. ∎
@@
As seen above, each $d_{k}$ is a retractive direction of $X$. However, the limit $(1,0)$ is not a retractive direction of $X$. To show this, we consider a sequence $\{\bar{x}_{k}\}\subset X$ given by
@@
Hence, $(1,0)$ is not a retractive direction of $X$, implying that $\mathcal{R}(X)$ is not closed. ∎
@@
Proposition 3.2.
Let $X=\cap_{i=1}^{m}X_{i}$ with closed sets $X_{i}$, for $m\geq 2$. If $X_{\infty}=\cap_{i=1}^{m}(X_{i})_{\infty}$, then $\cap_{i=1}^{m}\mathcal{R}(X_{i})\subseteq\mathcal{R}(X).$
@@
implying that $x_{k}-\rho d\in\cap_{i=1}^{m}X_{i}$ for all $k\geq K$. Thus, $d\in\mathcal{R}(X)$. ∎
When the sets $X_{i}$ in Proposition 3.2 are closed and convex, the condition $X_{\infty}=\cap_{i=1}^{m}(X_{i})_{\infty}$ is always satisfied, as seen from Proposition 2.1.
3.3 Cone of Retractive Directions of Functions
@@
By definition, one can show that $\mathcal{R}(f)$ is a cone and $0\in\mathcal{R}(f)$. Furthermore, for a proper function $f$ and a lower-level set $L_{\gamma}(f)$, we have that
@@
It turns out that, in general, there is no special relationship between the cone $\mathcal{R}(f)$ and the cone $\mathcal{R}(L_{\gamma}(f))$. The following example illustrates that we can have $\mathcal{R}(L_{\gamma}(f))\subseteq\mathcal{R}(f)$ for a nonconvex function.
@@
Let $f(s)=\sqrt{|s|}$. Then, for every $\gamma>0$, the lower-level set $L_{\gamma}(f)$ is nonempty and bounded, implying that $\mathcal{R}(L_{\gamma}(f))=\{0\}.$ However, we have that $({\rm epi}f)_{\infty}=\{(d,w)\mid d\in\mathbb{R},w\geq 0\}$, implying that $f_{\infty}(d)=0$ for all $d\in\mathbb{R}$. Moreover, in this case $\mathcal{R}(f)=\mathbb{R}$, and we have $\mathcal{R}(L_{\gamma}(f))=\{0\}\subset\mathcal{R}(f)$ for any $\gamma>0$. ∎
In the following proposition, we establish some properties of $\mathcal{R}(f)$ for a convex function. Interestingly, in this case $\mathcal{R}(f)\subseteq\mathcal{R}(L_{\gamma}(f))$ for a lower-level set $L_{\gamma}(f)$ (converse inclusion to that of Example 3.9).
Proposition 3.3.
Let $f$ be a proper closed convex function. Then, we have
  1. (a)
@@
  2. (b)
    $\mathcal{R}(f)\subseteq\mathcal{R}(L_{\gamma}(f))$ for any $\gamma\in\mathbb{R}$.
@@
(a) Let $d\in\mathcal{R}(f)$. Then, for any $\{x_{k}\}$ converging in direction $d$ and any $\rho>0$, there is a large enough $K$ such that $f(x_{k}-\rho d)\leq f(x_{k})$ for all $k\geq K.$ Let $y_{k}=x_{k}-\rho d$. Then, $\{y_{k}\}$ also converges in direction $d$, and the preceding relation implies that
@@
By Proposition 2.5.2 of [2] for a proper closed convex function $f$, we have that
@@
Since $\rho>0$ and $f_{\infty}$ is positively homogeneous by Proposition 2.5.1 of [2], it follows that $f_{\infty}(d)\geq 0,$ which combined with the fact that $\mathcal{R}(f)\subseteq\mathcal{K}(f)$, implies that $f_{\infty}(d)=0$.
(b) Let $L_{\gamma}(f)$ be nonempty.
To arrive at a contradiction, assume that there is $d\in\mathcal{R}(f)$ such that $d\notin\mathcal{R}(L_{\gamma}(f))$. Since $\mathcal{R}(f)\subseteq\mathcal{K}(f)$, it follows that $d\in\mathcal{K}(f)$. By Proposition 2.3, we have that $(L_{\gamma}(f))_{\infty}=\mathcal{K}(f)$, implying that $d\in(L_{\gamma}(f))_{\infty}$. Since $d\in(L_{\gamma}(f))_{\infty}$ but it is a non-retractive direction of the set $L_{\gamma}(f)$, there is a sequence $\{x_{k}\}\subset L_{\gamma}(f)$ converging in direction $d$ and some $\bar{\rho}>0$ such that
@@
Without loss of generality, we may assume that $f(x_{k}-\bar{\rho}d)\geq\gamma$ for all $k$, for otherwise we would choose a suitable subsequence of $\{x_{k}\}$. Since $\{x_{k}\}\subset L_{\gamma}(f)$, it follows that
@@
Thus, we have that $\{x_{k}\}$ converges in direction $d$ and the direction $d$ belongs to $\mathcal{K}(f)$, but $f(x_{k}-\bar{\rho}d)\geq f(x_{k})$ for all $k$. Hence, the direction $d$ is not retractive for the function $f$, i.e., $d\not\in\mathcal{R}(f)$, which is a contradiction. ∎
The inclusion in Proposition 3.3(b) can be strict, as seen in the following example.
@@
It can be seen that $f_{\infty}(d)=\langle c,d\rangle$. By Proposition 3.3, if a direction $d$ is retractive, then $\langle c,d\rangle=0$.
Thus, $\mathcal{R}(f)=\{d\mid\langle c,d\rangle=0\}\subset\mathcal{R}(L_{\gamma}(f))$ for all $\gamma\in\mathbb{R}$. ∎
@@
Consider the function $f(x_{1},x_{2})=e^{-\sqrt{x_{1}x_{2}}}$ for $x_{1}\geq 0$, $x_{2}\geq 0$. The lower-level set $L_{\gamma}(f)$ is nonempty for all $\gamma>0$. Consider $L_{1}(f)$ and $L_{\gamma}(f)$ with $\gamma>0$. For the set $L_{1}(f)$ we have $L_{1}(f)=\mathbb{R}^{2}_{+},$ which is a polyhedral set, so we have $\mathcal{R}(L_{1}(f))=\mathbb{R}^{2}_{+}$. For the set $L_{\gamma}(f)$ with $\gamma>0$, we have
@@
Finally, we consider polynomial functions. A polynomial $h:\mathbb{R}^{n}\rightarrow\mathbb{R}$ of order $p$ has the following representation [5]:
@@
Using the relation $h(x)=\sum_{i=0}^{p}\phi_{i}(x)$, for all $x\in\mathbb{R}^{n}$, and the alternative characterization of the asymptotic function, as given in Theorem 2.2, we have
$h_{\infty}(d)=\liminf_{\begin{subarray}{c}t\rightarrow\infty\\ d^{\prime}\rightarrow d\end{subarray}}\frac{h(td^{\prime})}{t}$ $=\liminf_{\begin{subarray}{c}t\rightarrow\infty\\ d^{\prime}\rightarrow d\end{subarray}}\left(\sum_{i=0}^{p}t^{-1}\phi_{i}(td^{\prime})\right)$
@@
where the last equality follows from $\phi_{0}/t\to 0$, as $t\to\infty$, and the fact that each $\phi_{i}$ is a continuous function. When $\mu(d)\geq 2$, if $\phi_{\mu(d)}(d)>0$, then $h_{\infty}(d)=+\infty$, while if $\phi_{\mu(d)}(d)<0$, then $h_{\infty}(d)=-\infty$. When $\mu(d)=1$, then we are left with $\phi_{1}(d)$. ∎
@@
By Lemma 3.1 we have that
@@
Let $d$ be such that $h_{\infty}(d)=0$. Then, for every $i=1,\ldots,p$, we have $\phi_{i}(-d)=(-1)^{i}\phi_{i}(d)$, implying that $h_{\infty}(-d)=0$. Hence, by Theorem 2.5.3 in [2] it follows that
@@
where the last inclusion in the preceding relation holds since every direction $d\in\mathcal{C}(h)$ is retractive. By Proposition 3.3(a), we have that $\mathcal{R}(h)\subseteq\{d\mid h_{\infty}(d)=0\}$, which implies that the equality holds throughout in (7). ∎